# Balanced Similarity with Auxiliary Prompts: Towards Alleviating Text-to-Image Retrieval Bias for CLIP in Zero-shot Learning

### v1

@@ 6.4 Ablation Study (Q3)
For the similarity scores generated for query text and auxiliary prompts with many image regions, we have explored several straightforward and common methods for normalization. These include direct normalization, Min-max normalization, and softmax operation. The direct normalization technique scales, mapping the values to the range of 0 to 1. Min-max normalization, also known as feature scaling, transforms data linearly to a specified range, typically [0, 1], making it suitable for various analyses and machine learning algorithms. We compared these three operations and ultimately observed that the softmax operation is superior.
6.4.4 Analysis on Constant Parameter $\alpha$
@@
Since our task is zero-shot, we strive to select constant parameters $\alpha$ using as small a dataset as possible. In formula 7 of the hybrid method, we use $\alpha=0.75$ for the REC task and $\alpha=0.5$ for the RIS task in our final experiments. We conducted parameter experiments on the validation set of the RefCOCOg dataset, as shown in Figure 8 and Figure 9. Notably, for the RIS task, we exclusively employ ViT-B/32 as the backbone for CLIP and evaluation metric mIoU to investigate the parameter $\alpha$. For the constant parameter $\alpha$, optimal values may exist for different tasks, datasets, CLIP backbones, and evaluation metrics.
Figure 8: In the REC task on the validation set of the RefCOCOg dataset, the impact of different $\alpha$ values on accuracy is examined through Reclip’s application of hybrid similarity.
Figure 9: In the RIS task on the validation set of the RefCOCOg dataset, the impact of different $\alpha$ values on mIoU is examined through Zero-shot-RIS’s application of hybrid similarity. We have selected ViT-B/32 as the backbone for CLIP.
6.5 Templates used for Auxiliary Prompts
Different templates can influence our auxiliary prompts, thus affecting the final results. In the RIS task, we investigated the impact of varying template lengths on the ultimate outcome. While changes in template length can affect the performance of our BASI, even without using a template, our BASI still outperforms the original similarity on the RIS task.
Below are the templates used in our experiments:
  • Length = 1
    “A <>absent<>< >.”
  • Length = 2
    “Snapshot of <>absent<>< >.”
    “Image capturing <>absent<>< >.”
    “Picture featuring <>absent<>< >.”
    “Photo showcasing <>absent<>< >.”
    “Shot of <>absent<>< >.”
    “It’s a <>absent<>< >.”
    “Frame capturing <>absent<>< >.”
    “Portrait featuring <>absent<>< >.”
    “Scene of <>absent<>< >.”
    “Visual of <>absent<>< >.”
  • Length = 3
    “Image capturing <>absent<>< > moment.”
    “Snapshot showcasing <>absent<>< > beauty.”
    “Picture portraying <>absent<>< > scene.”
    “Photograph capturing <>absent<>< > essence.”
    “Shot revealing <>absent<>< > subject.”
    “Visualizing <>absent<>< > through lens.”
    “Photo depicting <>absent<>< > atmosphere.”
    “Capture highlighting <>absent<>< > detail.”
    “Picture framing <>absent<>< > story.”
    “Image capturing <>absent<>< > essence.”
  • Length = 4
    “A snapshot capturing <>absent<>< > beautifully.”
    “An image showcasing <>absent<>< > magnificently.”
    “A picture framing <>absent<>< > perfectly.”
    “A photograph highlighting <>absent<>< > elegantly.”
    “A shot revealing <>absent<>< > vividly.”
    “A view capturing <>absent<>< > artistically.”
    “A photo featuring <>absent<>< > splendidly.”
    “A visual story of <>absent<>< >.”
    “A scene portraying <>absent<>< > wonderfully.”
    “A picture telling <>absent<>< > tale.”
  • Length = 5
    “Capturing the essence of <>absent<>< > beautifully.”
    “A snapshot freezing the moment of <>absent<>< >.”
    “An image portraying the beauty of <>absent<>< >.”
    “A picture capturing the essence of <>absent<>< >.”
    “A photograph showcasing the beauty of <>absent<>< >.”
    “A shot revealing the details of <>absent<>< >.”
    “A visual story of <>absent<>< > captured.”
    “A frame encapsulating the essence of <>absent<>< >.”
    “A portrait highlighting the beauty of <>absent<>< >.”
    “You can see <>absent<>< > in the picture.”
  • Length = 6
    “Captured a stunning photo of <>absent<>< > yesterday.”
    “Took a beautiful photo of <>absent<>< > during sunset.”
    “Snapped an unforgettable photo of <>absent<>< > downtown.”
    “Managed to get a perfect photo of <>absent<>< >.”
    “Caught on camera: an incredible photo of <>absent<>< >.”
    “Shared a mesmerizing photo of <>absent<>< > on social media.”
    “Recently framed a lovely photo of <>absent<>< > together.”
    “Documented the moment with a special photo of <>absent<>< >.”
    “Featured in the gallery: a breathtaking photo of <>absent<>< >.”
    “Just developed a cherished photo of <>absent<>< > today.”
  • Length = 7
    “Capturing the essence: a photo of <>absent<>< >.”
    “An exquisite composition: a photo capturing <>absent<>< >.”
    “In the frame, a stunning image of <>absent<>< >.”
    “A snapshot frozen in time: the beauty of <>absent<>< >.”
    “Witness the moment through this captivating photo of <>absent<>< >.”
    “An eight-word story told in a photo: <>absent<>< >.”
    “Through the lens: a mesmerizing portrayal of <>absent<>< >.”
    “A visual narrative unfolds in this photo of <>absent<>< >.”
    “Discover the details within this compelling photograph of <>absent<>< >.”
    “In focus: an evocative image telling <>absent<>< > story.”
7 Conclusion
In this paper, we first discovered the text-to-image retrieval bias of CLIP and figured out that the imbalanced similarity score ranges cause some of the bias. Then, we proposed a Balanced Similarity with Auxiliary Prompts (BASP), an effective plug-in module for CLIP to the text-to-image retrieval bias and boosted performance in zero-shot learning. Extensive experiments using REC and RIS with three datasets demonstrated the capability of our CLIP. There is still room for improvement in our BASP, such as generating appropriate prompts depending on each text query description. Besides, other cross-modal foundation models, including BLIP [21] and BLIP-v2 [24], may also contain the image-to-text retrieval bias. How to apply our BASP to other foundation models is valuable for future work. Furthermore, there may exist image-to-text retrieval bias in CLIP, and analyzing the image-to-text retrieval bias is also one of our future works.

### v2

@@ VI-D2 Effects of Template Length
For balanced similarity scores generated for query text and auxiliary prompts with many image regions, we have explored several straightforward and common methods for normalization. These include direct normalization, Min-max normalization, and softmax operation. The direct normalization technique scales, mapping the values to the range of 0 to 1. Min-max normalization, also known as feature scaling, transforms data linearly to a specified range, typically [0, 1], making it suitable for various analyses and machine learning algorithms. In Table VII, we compared these three operations and ultimately observed that the softmax operation is superior.
VI-D4 Analysis on Constant Parameter $\alpha$
@@
Since our task is zero-shot, we strive to select constant parameters $\alpha$ using as small a dataset as possible. In formula 7 of the hybrid method, we use $\alpha=0.75$ for the REC task and $\alpha=0.5$ for the RIS task in our final experiments. We conducted parameter experiments on the validation set of the RefCOCOg dataset, as shown in Fig. LABEL:fig:alpha (a) and Fig. LABEL:fig:alpha (b). Notably, for the RIS task, we exclusively employ ViT-B/32 as the backbone for CLIP and evaluation metric mIoU to investigate the parameter $\alpha$. For the constant parameter $\alpha$, optimal values may exist for different tasks, datasets, CLIP backbones, and evaluation metrics.
VI-E Applying BSAP to Different VLMs (Q4)
Figure 7: Apply BSAP to different VLMs in the text-to-image retrieval task using the CIFAR-10 dataset. The light-colored bars represent the number of correct predictions made directly by using VLM, while the dark-colored bars represent the number of correct predictions made by integrating BSAP into VLM for prediction.
We randomly sampled one image from each of the 10 classes of the CIFAR-10 dataset, using the category as our query text to see which image yielded the highest confidence score after being processed by the VLM. If the image corresponding to the highest confidence score was indeed from that category, the VLM’s judgment was considered correct. Conversely, if it was not from that category, the judgment was deemed incorrect. The experiment performed the aforementioned operation 100 times for each of the 10 categories, utilizing the CLIP, ALBEF, and BLIP models. As shown in Fig. 7, the light-colored bars indicate the number of successful judgments achieved. To demonstrate the reduction of hallucinations by our BSAP, we integrated BSAP into the three models and compared the balanced scores as shown by the dark-colored bars in Fig. 7. Here, BSAP refers to the application of other categories within CIFAR.
As can be observed from the bar chart, the performance of the models has been enhanced and hallucinations have been mitigated after the integration of BSAP. Improvements are noted across different VLMs.
We have also applied it to our two target tasks in Table I and Table II, REC and RIS. Since both ALBEF and BLIP were trained on the COCO dataset, they cannot be considered to be making zero-shot judgments. Nevertheless, the task still demonstrates the performance improvement of ALBEF and BLIP, indicating that our BSAP has a certain mitigating effect on the hallucination phenomenon in this class of large models.
VI-F Further Discussion of Applying BSAP in Image-to-Text Retrieval Task
Given that our primary focus is on hallucinations in zero-shot tasks for large models, specifically within the context of the text-to-image retrieval task, we have identified that BSAP can alleviate such hallucinations. Extending our investigation to image-to-text retrieval tasks, we conducted preliminary experiments. We randomly selected one image from each category in the CIFAR-10 dataset as our query image and tasked it with identifying the corresponding CIFAR-10 category. A correct identification of the category was considered a correct judgment; otherwise, it was deemed incorrect. Using both the original CLIP and the CLIP integrated with BSAP, we performed 100 draws for each of the ten categories, as detailed in Table VIII. For BSAP, an auxiliary prompt image was used, which was randomly selected from another category to serve as the BSAP. The experimental results indicate that, for the majority of categories, the use of BSAP provided significant guidance, suggesting that our BSAP also has a mitigating effect on hallucinations within image-to-text retrieval tasks.
VII Conclusion
In this paper, we first analyzed the text-to-image retrieval hallucination of CLIP and figured out that the imbalanced similarity score ranges cause some of the hallucinations. Then, we proposed Balanced Scores with Auxiliary Prompts (BSAP), an effective plug-in module for CLIP to the text-to-image retrieval hallucination, and boosted performance in zero-shot learning. Extensive experiments concerning REC and RIS with three datasets demonstrated the capability of our BRSP. There is still room for improvement in our BSAP, such as generating appropriate prompts depending on each text query description. Besides, other cross-modal foundation models, including ALBEF [45] and BLIP [46], may also contain the image-to-text retrieval hallucination. How to apply our BRSP to other foundation models is valuable for future work. Furthermore, there also exists the image-to-text retrieval hallucination in VLMs, and analyzing the image-to-text retrieval hallucination is also one of our future works.